OpenStudy (anonymous):

Find h'(1) given h(x) = f(x^(1/2)) f(1) = 9 f'(1) = 3

OpenStudy (anonymous):

well using chain rule h'(x) = d(x^(1/2))/dx * f'(x^(1/2)) h'(x) = (1/2)*x^(-1/2)*f'(x^(1/2)) sub in x = 1 h'(1) = (1/2)*1*f'(1) f'(1) = 3 so h'(1) = 3/2

OpenStudy (anonymous):

hope that helps, i'm not sure it's right because I didn't use f(1) = 9

OpenStudy (anonymous):

Nope, that was right. Thanks a lot! (the f(1) = 9 was for a different part of the problem, I figured I'd include it just in case I had been approaching the whole thing all wrong).

OpenStudy (anonymous):

lol